OpenStudy (prakharluv):

Weight of 1 atom of an element is 6.644*10^-23 gm. What are gm-atoms(moles) of element in 40kg?

thomaster (thomaster):

Do you mean how many atoms are in 40Kg of this element?

OpenStudy (prakharluv):

no....how many moles...

OpenStudy (chmvijay):

6.6410^-23 gm =1 atom 40 kg =40/6.6410^-23 = X find X 6.023 *10^23 atoms1 = 1 mole X atoms = X/ 6.023*10^23 =Y moles